# CHARTOM: A Visual Theory-of-Mind Benchmark for Multimodal Large Language Models

### v2

@@ Abstract
We introduce CHARTOM, a visual theory-of-mind benchmark for multimodal large language models. CHARTOM consists of specially designed data visualizing charts. Given a chart, a language model needs to not only correctly comprehend the chart (the FACT question) but also judge if the chart will be misleading to a human reader (the MIND question). Both questions have significant societal benefits. We detail the construction of the CHARTOM benchmark including its calibration on human performance. We benchmarked leading LLMs as of late 2024—including GPT, Claude, Gemini, Qwen, Llama, and Llava—on the CHARTOM dataset and found that our benchmark was challenging to all of them, suggesting room for future large language models to improve.
@@ 1 Introduction
For AI to better assist humans, it must know not just factual truth but also how humans perceive the truth. The two could differ for many reasons such as information asymmetry or cognitive ability. If AI can detect the difference, it can help humans accordingly. The ability of an individual to reason about how others think (rather than the factual truth) is known as the theory of mind [34]. A classic example is the Sally-Anne test [43, 3]: Sally hides a marble in a basket and then leaves the room. Anne moves the marble to a box while Sally is away. Sally comes back. At this point, one can ask an observer two types of questions:
@@
There is a recent surge of interest in AI theory of mind, where large language models take the place of the observer. Several researchers found current AI competent at theory of mind tasks, performing near or surpassing human performance [22, 39], though some others remained skeptical [41, 37]. Concerned with limitations of AI theory of mind benchmarks, recent work also started to develop new benchmarks emphasizing conversations [20], causality [12], and actions [47]. However, these existing AI theory of mind benchmarks are all based on text comprehension.
Our theory of mind benchmark tests visual perception instead. Specifically, our focus is on misleading data visualizing charts, e.g. bar charts, line graphs, pie charts, etc. Charts are widely used for conveying quantitative information to human readers. They lend credibility to accompanying messages and hence have a strong impact on human decision-making. On the flip side, however, if charts are misleading, they can amplify the impact of misinformation [24, 33]. Figure 1 reproduces a misleading chart whose $y$-axis is inverted, which may give the impression that enacting Florida’s “stand your ground” law reduced gun deaths. Misleading charts are unfortunately prevalent in many areas of our lives. For example, a review of medical advertisements found that 1/3 of charts were misleading [8]. A review of charts in Korean news about the COVID-19 pandemic showed that about 30% of bar charts and 44% of pictorial charts were misleading [23]. A U.S.-based analysis of visuals that had been identified as misleading by human fact-checkers showed that more than half of these visuals were used as evidence for fake news about COVID-19 [5]. Because journalists often use charts to make narratives appear objective and to engage viewers emotionally [36], misleading charts can make an inaccurate story more deceptive [9, 42].
Figure 1: Reproduction of a misleading chart from [11], attributed to Reuters originally
@@
1.1 Related Works
Existing computer vision research in charts have largely focused on chart comprehension. Classic approaches often employ an intermediate step of chart-to-data parsing, then use the data for downstream tasks such as visual question-answering, reasoning, and summarization. To this end, [6, 31] applied deep learning-based object detection methods to extract data, while others [1, 7, 29, 18, 38] applied semi-automated off-the-self OCR text recognition systems [4, 19]. Newer LLM-based approaches tend to be end-to-end, such as training the model on extensive pre-training datasets [28, 14, 30, 46], few-shot tuning methods which fine-tune pre-trained models with a small number of task-specific examples [10, 26, 21], and even zero-shot learning where the model makes predictions on tasks it has not been explicitly trained on [44, 17]. The newer approaches are performant, though occasionally still suffer from hallucination in data extraction and reasoning [16, 15].
However, we point out a significant shortcoming in chart comprehension: many misleading charts in fact contain the correct data. Take Figure 1 for example. Despite the unusual flipped $y$-axis, a meticulous reader (or AI) can indeed recover all the correct (year, deaths) data points from the chart. Therefore, ironically, a perfect chart comprehension system is not directly suited for judging how misleading a chart may be to humans. This is the same distinction between factual and theory of mind questions. We can ask:
@@ 1.1 Related Works
Current chart comprehension systems aim to solve the FACT question. To the best of our knowledge, the present paper is the first to address the MIND question in charts instead.
Our Contributions:
Towards this end, we make the following contributions:
@@
    We create a visual theory-of-mind benchmark CHARTOM (for CHARt Theory Of Mind) consisting of specially designed charts. The benchmark covers common variety of chart types and planted visual manipulations in those charts as shown in Table 1
@@
    We designed two types of questions for each chart: 1.) FACT question that are based on factual information present in charts, and, 2.) MIND question based on human theory of mind.
@@
    The key to the benchmark is to obtain MIND ground truth, i.e. “degree of misleadingness to humans”, for each chart that has been obtained via human experiments. This ground truth enables accessing whether an AI correctly answers the MIND questions.
@@
    We test the performance of leading LLMs in Claude, Gemini, GPT, Llama, LLava, and Qwen series as of Nov 2024 and report our findings. We observe that all LLMs perform poorly on both FACT and MIND questions indicating a big scope of improvement.
@@ 2 The CHARTOM Benchmark
Figure 2: A pair of non-manipulated(left) and manipulated(right) charts in CHARTOM with truncated y-axis manipulation.
@@ 2.1 Design of the CHARTOM Benchmark
An example. The pair of charts in Figure 2 share the same underlying data: (year=2018, dogs=250) and (year=2019, dogs=300). The left chart has a standard design which we call “non-manipulated”, while in the right chart we intentionally planted a potentially misleading manipulation: we truncated the $y$-axis range so it starts at 240 instead of 0. This is to create a perceptual effect: the 2019 bar now appears much taller than the 2018 bar. For each of the two charts, we ask the following FACT and MIND questions.
@@
The 112 charts come in 56 pairs (one pair is already shown in Figure 2). Each pair has a non-manipulated (suffix _1) and a manipulated (suffix _2) version, the latter incorporates one of many visualization fallacies suggested by the psychology literature, see the ‘manipulation’ column in Table 1. The manipulated version is intended to mislead casual human readers. Each pair of charts also comes with a FACT question and a MIND question. The FACT question is typically a simple word problem based on chart comprehension, see examples in Appendix A. There are three types of FACT questions: multiple-choice, free text entry, and sorting. Importantly, the true underlying data can still be recovered from either version if one is careful. This means that the FACT question has the same answer on both versions. The MIND question has the same format as shown earlier for all charts.
To summarize, the CHARTOM benchmark consists of
  • 112 charts: 56 pairs of non-manipulated and manipulated versions.
  • Each chart comes with a FACT question and the answer key.
  • Each chart comes with a MIND question, and the answer key is the Human Misleadingness Index (HMI, see Section 3.2) observed in human experiments.
To avoid web crawlers inadvertently adding the benchmark to future LLM training data, the CHARTOM benchmark is not online but is available upon request.
Non-Manipulated Manipulated
chart type group ID HMI manipulation type HMI
line G1 0.51 inconsistent x-axis 0.68
line G2 0.76 compressed y-axis 0.88
line G3 0.03 inverted y-axis 0.83
line G4 0.05 truncated y-axis 0.92
line G9 0.19 dual axes 0.84
bar G7 0.31 truncated y-axis 0.53
bar G8 0.76 compressed y-axis 0.69
bar G10 0.19 pictorial bars 0.27
bar G12 0.11 3D effect 0.48
scatter G5 0.02 logarithmic y-axis 0.63
scatter G6 0.02 swapped axis 0.73
map G15 0.02 inverted color scale 0.15
pie G13 0.25 3D effect 0.79
pie G14 0.30 3D effect + pop out 0.79
@@ 3.1 Ground Truth for FACT Questions
    Numerical (free text) questions. For numerical answers, the answer key provides the exact answer assuming perfect perception. But an exact match can be too strong an assumption for either humans or LLMs when they can enter any free-text number. To judge numerical answers, we suggest using a ±$\pm 10\%$ tolerance interval: if the answer key is $a$, then an LLM answer in the interval $[0.9a,1.1a]$ is deemed correct.
@@ 3.2 Ground Truth for MIND Questions
In contrast, the MIND question is trickier to judge. On the non-misleading chart in Figure 2(left) where the design is standard, one might still argue that the chart is not perfect and could be a little bit misleading to humans: the numbers are too small, there is no grid guidelines to help visual alignment, etc. Conversely, on the misleading chart (right) even if we intentionally planted the truncated $y$-axis because we presumed it to be misleading based on existing psychology literature (e.g. [27]), we do not know a priori the actual effect on human readers. To avoid these difficulties we propose the Human Misleadingness Index (HMI), a number between 0 and 1 for each chart which is the fraction of the human population that is misled by the chart. The correct answer to a MIND question is the HMI value.
To obtain HMI we conducted human experiments, which is expensive and time consuming but is the only proper method. Our human experiment involves 68 university students with IRB approval. We did not directly ask human participants the MIND question because we do not anticipate people to be able to give reliable answers—after all, many were themselves misled by some of our charts. Instead, we asked them the FACT question, which is designed so that users who are misled by the chart are likely to give wrong answers. We randomized the order of non-manipulated and manipulated charts in each pair when presenting them to human participants to minimize the order effect. We estimate the HMI of a chart by the percent of our study population who did not give a correct answer on the FACT question.
@@
Figure 3 shows the HMI values in the CHARTOM benchmark. These values are sorted from small (not misleading to humans) to large (very misleading). In addition, each HMI value is color-coded to indicate whether the chart is the non-manipulated version (blue) or the manipulated version (red). Generally speaking, the non-manipulated charts have small HMI, and the manipulated ones have large HMI. But this depends on the type of manipulation. As the HMI values in Table 1 indicate, some manipulations greatly misled people: inconsistent x-axis, inverted y-axis, dual axes plots, 3D effect, etc. While interestingly, some other manipulations did not confuse people much: compressed y-axis and pictorial bars are benign. Such differences are of interest to cognitive science. Our human experiments are described in greater detail in [35].
@@ 4 Benchmarking Leading LLMs
We evaluated leading LLMs including Claude Sonnet 3.5 [2], Gemini 1.5 Pro [40], GPT-4o [32], LLaVA-OneVision [25], Llama-3.2-90B [13], Qwen2-VL-72B [45] on both FACT and MIND questions and present our findings below. When evaluating LLMs, we suggest the following criteria:
  • On FACT questions, we suggest reporting accuracy.
@@
4.1 Like Humans, LLMs are Fooled by Manipulations
We asked LLMs to answer FACT questions on both non-manipulated and manipulated charts and record their response. Figure 4 shows the error rate of each LLM separately on non-manipulated charts (blue), manipulated charts (orange), and all charts (green). We make the following observations:
@@ 4.1 Like Humans, LLMs are Fooled by Manipulations
    LLMs perform significantly worse on manipulated charts as compared to non-manipulated ones. This is evident from the difference in height of blue and orange bars for each LLM in Figure 4. We performed a $\mathcal{X}^{2}$ test that achieved a p-value $\leq 0.05$ for all LLMs. This suggests that, like humans, LLMs are fooled by manipulations.
@@
    Top LLMs show near or superhuman performance on FACT questions. We observe that top LLMs like Claude, Qwen, and GPT-4o have performed better or very close to humans (on average) while others like Llama, Gemini, and LLaVA have shown relatively poor performance.
  3. 3.
    Some LLMs fail to understand factual information in non-manipulated charts. While LLMs like Claude, Qwen, and GPT-4o excel on non-manipulated charts, they still make mistakes on 15-25% of cases. In contrast, models like Llama, Gemini, and LLaVA perform significantly worse indicating a big scope of improvement even on clean charts.
Figure 4: The error rate of humans and LLMs on FACT questions

### v3

# CHARTOM: A Visual Theory-of-Mind Benchmark for LLMs on Misleading Charts

@@ Abstract
We introduce CHARTOM, a visual theory-of-mind benchmark designed to evaluate multimodal large language models’ capability to understand and reason about misleading data visualizations though charts. CHARTOM consists of carefully designed charts and associated questions that require a language model to not only correctly comprehend the factual content in the chart (the FACT question) but also judge whether the chart will be misleading to a human readers (the MIND question), a dual capability with significant societal benefits. We detail the construction of our benchmark including its calibration on human performance and estimation of MIND ground truth called the Human Misleadingness Index. We evaluated several leading LLMs—including GPT, Claude, Gemini, Qwen, Llama, and Llava series models—on the CHARTOM dataset and found that it was challenging to all models both on FACT and MIND questions. This highlights the limitations of current LLMs and presents significant opportunity for future LLMs to improve on understanding misleading charts.
1 University of Wisconsin-Madison   2 ETH Zurich   3 Google DeepMind
@@ 1 Introduction
For AI to better assist humans, it must know not just factual truth but also how humans perceive the truth. The two could differ for many reasons such as information asymmetry or cognitive ability. If AI can detect the difference, it can help humans accordingly. The ability of an individual to reason about how others think (rather than the factual truth) is known as the theory of mind premack1978does . A classic example is the Sally-Anne test wimmer1983beliefs ; baron1985does : Sally hides a marble in a basket and then leaves the room. Anne moves the marble to a box while Sally is away. Sally comes back. At this point, one can ask an observer two types of questions:
@@
There is a recent surge of interest in AI theory of mind, where large language models take the place of the observer. Several researchers found current AI competent at theory of mind tasks, performing near or surpassing human performance kosinskitheory ; strachan2024testing , though some others remained skeptical ullman2023large ; sap2022neural . Concerned with limitations of AI theory of mind benchmarks, recent work also started to develop new benchmarks emphasizing conversations kim2023fantom , causality gandhi2024understanding , and actions zhou2023far . However, these existing AI theory of mind benchmarks are all based on text comprehension.
Our theory of mind benchmark tests visual perception instead. Specifically, our focus is on misleading data visualizing charts, e.g. bar charts, line graphs, pie charts, etc. Charts are widely used for conveying quantitative information to human readers. They lend credibility to accompanying messages and hence have a strong impact on human decision-making. On the flip side, however, if charts are misleading, they can amplify the impact of misinformation lauer2020people ; pandey2015deceptive . Figure 1(a) reproduces a misleading chart whose $y$-axis is inverted, which may give the impression that enacting Florida’s “stand your ground” law reduced gun deaths. Misleading charts are unfortunately prevalent in many areas of our lives. For example, a review of medical advertisements found that 1/3 of charts were misleading cooper2003quantity . A review of charts in Korean news about the COVID-19 pandemic showed that about 30% of bar charts and 44% of pictorial charts were misleading kwon2021graphs . A U.S.-based analysis of visuals that had been identified as misleading by human fact-checkers showed that more than half of these visuals were used as evidence for fake news about COVID-19 brennen2021beyond . Because journalists often use charts to make narratives appear objective and to engage viewers emotionally ryan2016visual , misleading charts can make an inaccurate story more deceptive diakopoulos2018ethics ; west2021misinformation .
(a) Reproduction of a misleading chart from BusinessInsider , attributed to Reuters originally
(b) Non-manipulated bar chart
(c) Manipulated bar chart
A pair of example charts in CHARTOM
Figure 1: Example of manipulated charts in the wild and in our CHARTOM dataset.
@@
However, we point out a significant shortcoming in chart comprehension: many misleading charts in fact contain the correct data. Take Figure 1(a) for example. Despite the unusual flipped $y$-axis, a meticulous reader (or AI) can indeed recover all the correct (year, deaths) data points from the chart. Therefore, ironically, a perfect chart comprehension system is not directly suited for judging how misleading a chart may be to humans. This is the same distinction between factual and theory of mind questions. We can ask:
@@
Current chart comprehension systems aim to solve FACT questions. To the best of our knowledge, the present paper is the first to address the MIND question in charts instead.
Our Contributions: Towards this end, we make the following contributions:
@@
    We create a visual theory-of-mind benchmark CHARTOM (for CHARt Theory Of Mind) consisting of specially designed charts. The benchmark covers common variety of chart types including line, bar, scatter, map, and pie charts as shown in Table 1
@@
    We designed two types of questions for each chart: 1.) a FACT question that tests the understanding of factual information present in charts, and, 2.) a MIND question that tests the understanding of “how much” and “why” a chart is misleading to humans.
@@
    The key to the benchmark is MIND ground truth called HMI (Section 4.2) that quantifies the “degree of misleadingness to humans”, of each chart that has been obtained via human experiments. We also provide ground truth reasoning that captures why humans are misled.
@@
    We test the performance of some of the leading LLMs in Claude, Gemini, GPT, Llama, Llava, and Qwen series and report our findings in Section 5. We observe that most LLMs perform poorly not only on MIND but also on FACT questions indicating their failure to understand manipulated charts properly. This also suggests a big scope for improvement in future LLMs to understand misleading charts.
2 Related Works
Recent works have extensively explored the capability of modern large language models (LLMs) to understand and reason about theory of mind(TOM) problems, focusing on narrative or conversational tasks involving belief attribution and social cognition kosinskitheory ; street2024motom ; chen2024tombench . Prior standard TOM benchmarks typically consist of tasks such as false-belief reasoning kosinskitheory , second-order inference street2024motom , and social reasoning tasks like faux pas detection, hint interpretation, irony, and sarcasm understanding strachan2024human all based on textual data. These studies have demonstrated that advanced LLMs, particularly GPT-4, approach or even surpass human-level performance on various TOM tasks chen2024tombench ; strachan2024human . However, despite the promising results, LLMs often exhibit brittleness, and struggle with implicitly conveyed mental states shapira2023faux ; chen2024tombench . Unlike these works, we focus on a very different visual theory of mind problem for understanding charts that mislead humans into drawing the wrong conclusion about presented data.
Existing computer vision research in charts has largely focused on chart comprehension. Classic approaches often employ an intermediate step of chart-to-data parsing, then use the data for downstream tasks such as visual question-answering, reasoning, and summarization. To this end, visual2019 ; chartvi2022 applied deep learning-based object detection methods to extract data, while others bar2015 ; scatteract2017 ; ivolver2016 ; chartsense2017 ; revision2011 applied semi-automated off-the-self OCR text recognition systems datathief ; tesseract2007 . Newer LLM-based approaches tend to be end-to-end, such as training the model on extensive pre-training datasets unichart2023 ; chartllama2023 ; chartassisstant2024 ; zhang2024tinychart , few-shot tuning methods which fine-tune pre-trained models with a small number of task-specific examples do2023llms ; liu2023deplot ; kim2024simplot , and even zero-shot learning where the model makes predictions on tasks it has not been explicitly trained on wu2024evaluating ; islam2024large . The newer approaches are performant, though occasionally still suffer from hallucination in data extraction and reasoning huang2024pixels ; advances2024 . Nevertheless, all the prior works have mainly focused on standard chart comprehension which is very different from our theory of mind understanding of misleading charts.
@@ 3.1 Design of the CHARTOM Benchmark
An example. The pair of charts in Figure 1 share the same underlying data: (year=2018, dogs=250) and (year=2019, dogs=300). The left chart has a standard design which we call “non-manipulated”, while in the right chart, we intentionally planted a potentially misleading manipulation: we truncated the $y$-axis range so it starts at 240 instead of 0. This is to create a perceptual effect: the 2019 bar now appears much taller than the 2018 bar. For each of the two charts, we ask the following FACT and MIND questions.
@@
The 112 charts come in 56 pairs (one pair is already shown in Figure 1). Each pair has a non-manipulated (suffix _1) and a manipulated (suffix _2) version, the latter incorporates one of many visualization fallacies suggested by the psychology literature, see the ‘manipulation’ column in Table 1. The manipulated version is intended to mislead casual human readers. Each pair of charts also comes with a FACT question and a MIND question. The FACT question is typically a simple word problem based on chart comprehension, see examples in Appendix A. There are three types of FACT questions: multiple-choice, free text entry, and ranking. Importantly, the true underlying data can still be recovered from either version if one is careful. This means that the FACT question has the same answer on both versions. The MIND question has the same format as shown earlier for all charts.
To summarize, the CHARTOM benchmark consists of:
  • 112 charts: 56 pairs of non-manipulated and manipulated charts.
  • Each chart comes with a FACT question and an answer key.
  • Each chart comes with a MIND question, and the answer key is the Human Misleadingness Index (HMI, see Section 4.2) obtained from human experiments.
chart type answer type group ID HMI(non-manip) manipulation type HMI(manip)
line free text G1 0.51 truncated y-axis 0.68
line free text G2 0.76 compressed y-axis 0.88
line multiple choice G3 0.03 inverted y-axis 0.83
line multiple choice G4 0.05 inconsistent x-axis 0.92
line multiple choice G9 0.19 dual axes 0.84
bar free text G7 0.31 truncated y-axis 0.53
bar free text G8 0.76 compressed y-axis 0.69
bar free text G10 0.19 pictorial bars 0.27
bar multiple choice G12 0.11 3D effect 0.48
scatter multiple choice G5 0.02 logarithmic y-axis 0.63
scatter multiple choice G6 0.02 inverted x-axis 0.73
map multiple choice G15 0.02 inverted color scale 0.15
pie rank G13 0.25 3D effect 0.79
pie rank G14 0.30 3D effect + pop out 0.79
@@ 4.1 Ground Truth for FACT Questions
    Numerical (free text) questions. We provide a correct answer key assuming perfect perception. But an exact match can be too strong an assumption for either humans or LLMs when they can enter any free-text number. So, we suggest using a ±$\pm 10\%$ tolerance interval: if the answer key is $a$, then an answer in the interval $[0.9a,1.1a]$ is deemed correct.
@@ 4.2 Ground Truth for MIND Questions
In contrast, the MIND question is trickier to judge. On the non-misleading chart in Figure 1(left) where the design is standard, one might still argue that the chart is not perfect and could be a little bit misleading to humans: the numbers are too small, there is no grid guidelines to help visual alignment, etc. Conversely, on the misleading chart (right) even if we intentionally planted the truncated $y$-axis because we presumed it to be misleading based on existing psychology literature (e.g. lo2022misinformed ), we do not know a priori the actual effect on human readers. To avoid these difficulties we propose a metric called Human Misleadingness Index as defined next.
Human Misleadingness Index (HMI) of a chart is a number between 0 and 1 that quantifies the fraction of the human population that is misled by the chart.
The correct answer to a MIND question is the HMI value. To obtain HMI we conducted human experiments, which is expensive and time-consuming but is the only proper method. Our human experiment involves 68 university students with IRB approval. We did not directly ask human participants the MIND question because we do not anticipate people to be able to give reliable answers—after all, many were themselves misled by some of our charts. Instead, we asked them the FACT question, which is designed so that users who are misled by the chart are likely to give wrong answers. We randomized the order of non-manipulated and manipulated charts in each pair when presenting them to human participants to minimize the order effect. We estimate the HMI of a chart by the percent of our study population who did not give a correct answer on the FACT question.
@@
Figure 2 shows the HMI values in the CHARTOM benchmark. These values are sorted from small (not misleading to humans) to large (very misleading). In addition, each HMI value is color-coded to indicate whether the chart is the non-manipulated version (blue) or the manipulated version (orange). Generally speaking, the non-manipulated charts have small HMI, and the manipulated ones have large HMI. But this depends on the type of manipulation. As the HMI values in Table 1 indicate, some manipulations greatly misled people: inconsistent x-axis, inverted y-axis, dual axes plots, 3D effect, etc. While interestingly, some other manipulations did not confuse people much: compressed y-axis and pictorial bars are benign. Such differences are of interest to cognitive science. Important details of our human experiments are presented in the appendix.
@@ 5 Benchmarking Leading LLMs
We evaluated leading LLMs including Claude-3.7-Sonnet (claude-3.7) anthropic2025sonnet3.7 , Claude-3.5-Sonnet (claude-3.5) anthropic2024Claude3family , Gemini-1.5-Pro (gemini-1.5) Geminiteam2024Gemini15unlockingmultimodal , Gemini-2.5-Flash (gemini-2.5) google2025gemini2.5 , GPT-4o (gpt-4o) openai2024gpt4ocard , OpenAI o1 (gpt-o1) openai2024openaio1card , Llama-3.2-90B (llama-3) grattafiori2024Llama3herdmodels , LLaVA-OneVision-72B (llava) li2024LLaVAonevisioneasyvisualtask , Qwen2-VL-72B (qwen-2) yang2024Qwen2technicalreport , Qwen-2.5-VL-72B (qwen-2.5) bai2025qwen25vltechnicalreport on both FACT and MIND questions and present our findings below. When evaluating LLMs, we suggest the following criteria:
  • On FACT questions, we suggest reporting accuracy based on the evaluation in section 4.1.
@@
Figure 3: The error rate of human participants and LLMs on FACT questions
@@
5.1 Like Humans, all LLMs are Fooled by Manipulations
We asked LLMs to answer FACT questions on both non-manipulated and manipulated charts and record their response. We then evaluate their performance based on the metric suggested above. Figure 3 shows the error rate of each LLM separately on non-manipulated charts (blue), manipulated charts (orange), and all charts (green). We make the following observations:
@@ 5.1 Like Humans, all LLMs are Fooled by Manipulations
    Top LLMs show near or superhuman performance on FACT questions. We observe that top LLMs like claude-3.7, gemini-2.5, qwen-2.5, and gpt-4o performed better or very close to humans (on average) while others like llama-3, gemini-1.5, and llava have shown relatively poorer performance.
@@
    LLMs perform significantly worse on manipulated charts as compared to non-manipulated ones. This is evident from the difference in height of blue and orange bars for each LLM in Figure 3. We performed a $\mathcal{X}^{2}$ test that achieved a p-value $\leq 0.05$ for all LLMs. This suggests that, like humans, LLMs are fooled by manipulations and there is big scope for improvements, particularly on manipulated charts.
    Furthermore, Some LLMs like qwen-2.5, gpt-o1, llama-3, gpt-4o fail to even understand non-manipulated charts. This suggests a good scope for future improvements.